# RealCustom++: Representing Images as Real-Word for Real-Time Customization

### v1

@@ 3.3.2 Generation Branch
In the GenB, the influence scope $\boldsymbol{\hat{M}}$ is multiplied by the visual cross-attention results to mitigate any negative impacts on the controllability of the given texts in other subject-irrelevant regions. Specifically, Eq. 4 is reformulated as:
@@
where the necessary resize operation is applied to match the size of $\boldsymbol{\hat{M}}$ with the resolution of each cross-attention block. The denoised output of GenB is denoted as $\boldsymbol{z_{t-1}}$. Classifier-free guidance [85] is applied to produce the next step’s denoised latent feature $\boldsymbol{z_{t-1}}$ as:
$\boldsymbol{z_{t-1}}=\epsilon_{\theta}(\emptyset)+\omega(\boldsymbol{z_{t-1}}-\epsilon_{\theta}(\emptyset)),$ (17)
where $\epsilon_{\theta}(\emptyset)$ is the unconditional denoised output and $\omega$ is the classifier-free guidance strength.
Figure 8: Illustration of the motivation behind the early stop regularization for mask calculation, where the image guidance mask tends to converge in the middle diffusion steps and become more scattered in the later steps.
@@ 3.4 Extension To Multi-subjects Customization
In contrast to the existing pseudo-word paradigm, which learns the correspondence between each pseudo-word and its respective subject, RealCustom++ establishes a general alignment between vision conditions and all real words in the text during training. This allows users to freely select different real words to customize various subjects during inference. This inherent flexibility makes RealCustom++ readily extendable to multiple-subject customization.
Extension to decouple multiple subjects from a single reference image (One2Many). Unlike previous foreign pseudo-word paradigm methods that require region-wise attention regularization or pre-defined masks to separate multiple subjects within a single reference, RealCustom++ can accomplish this task without any additional design: simply utilize different native real words to represent the target subjects in the reference images. As illustrated in Figure 2(b), using ’boy’ to customize only the target boy with green hair or simultaneously using ’boy’ and ’horse’ to customize both subjects, the disentanglement is naturally achieved thanks to the well-learned alignment between vision conditions and original text conditions.
Extension to compose multiple subjects from multiple reference images (Many2Many). Given $N$ reference images, each representing a different subject, we extend the Top-K selection to the Multi-Subject Top-K selection as illustrated in Alg.1. The key is to iteratively select the maximum scores for each given subject while ensuring that these guidance masks do not overlap. After acquiring the selected attention map for each subject $\boldsymbol{\bar{M}}_{j},j\in[0,N-1]$, maximum normalization is applied to each:
$\boldsymbol{\hat{M}}_{j}=\frac{\boldsymbol{\bar{M}}_{j}}{\text{max}(\boldsymbol{\bar{M})}_{j}},j\in[0,N-1].$ (18)
@@
TABLE I: Quantitative comparisons. RealCustom++ outperforms existing methods across all metrics on both SD-v1.5 and SDXL. Specifically: (1) For text controllability, on the more powerful base model SDXL, we achieve improvements of 7.01%, 7.48%, and 76.43% on CLIP-B-T, CLIP-L-T, and ImageReward, respectively. The significant improvement on ImageReward validates that RealCustom++’s non-conflict subject representation and the disentangled influence scope between texts and subjects can generate customized images with much higher quality (higher aesthetic score), as other subject-irrelevant regions are completely controlled by the given text without contamination from the subjects. (2) For subject similarity, we achieve state-of-the-art performance on CLIP-B-I, CLIP-L-I, and DINO-I, validating that RealCustom++ can generate more similar subjects in the subject-relevant regions.
4 Experiments
4.1 Experimental Setups
Implementation Details. RealCustom++ is implemented based on the pretrained Stable Diffusion 1.5[8] and Stable Diffusion XL[9] to further validate the generalization. We use a mix of the generic <text-image> dataset (i.e., the filtered subset of Laion-5B [83] based on aesthetic score) and the <multi-view> dataset (i.e., MVImageNet[82]) as detailed in Sec.3.2.2. RealCustom++ is trained using 8 A100 GPUs for 160,000 iterations with a learning rate of 1e-4. For Stable Diffusion XL, we first train the model at a 512 resolution for 60,000 iterations and then further fine-tune it at a 1024 resolution for another 100,000 iterations. We use SigLip[81] and DINO[80] with a $384^{2}$ input image size as our image encoders, concatenating their last hidden states in the channel dimension to formulate the deep image features $\in\mathbb{R}^{729\times 2176}$. Therefore, the notation $n_{\text{image}}=729$ and $c_{\text{image}}=2176$. The shallow layer number $L=3$, and the selected shallow layers for SigLIP and DINO are {7, 13, 19} and {4, 10, 16}, respectively. These layers are also concatenated in the channel dimension correspondingly, resulting in triple-level shallow image features. Unless otherwise specified, for Stable Diffusion 1.5, the DDIM sampler [57] with 50 sample steps is used for sampling, and the classifier-free guidance $\omega$ is set to 12.5. For Stable Diffusion XL, the DDIM sampler [57] with 25 sample steps is used for sampling, and the classifier-free guidance $\omega$ is set to 7.5. The Top-K ratio $\gamma_{\text{scope}}$ is set to 0.2 by default. The timestep threshold $T_{\text{stop}}$ for mask calculation early stop regularization is set to 25 for SD-v1.5 and 12 for SDXL.
(1) Subject-Similarity: We use the state-of-the-art segmentation model, SAM [92], to segment the subject, and then evaluate with both CLIP-I and DINO [80] scores. These scores represent the average pairwise cosine similarity of CLIP or DINO embeddings of the segmented subjects in generated and given images. For a more robust evaluation, we adopt two variants of CLIP: CLIP ViT-B/32 (denoted as CLIP-B-I) and CLIP ViT-L/14 (denoted as CLIP-L-I).
Evaluation Benchmarks. Following previous works, we use the prompt ”a photo of [category]” to generate images for calculating similarity. The full editing prompts and subject images used for evaluation are provided in the Appendix. These editing prompts and subject images are derived from the standard DreamBench[3] and are further supplemented with more challenging cases in open domains, such as cartoon characters and buildings, for a more comprehensive evaluation than our conference version.
Figure 9: Qualitative comparison with existing state-of-the-art customization methods. RealCustom++ produces higher quality results with better subject similarity and text controllability compared to existing methods. Additionally, RealCustom++ demonstrates superior subject diversity in poses and expressions (e.g., ”smile on her face” in the first row). Furthermore, RealCustom++ generates customization results with cleaner and more diverse backgrounds. For instance, in the second row, existing methods like Custom-Diffusion and MS-Diffusion inherit the pink background from the given subject image, while RealCustom++ generates a background that aligns more closely with the prompt ”on the beach.”
Prior SOTAs. We compare with existing state-of-the-art text-to-image customization methods includes both optimization-based and optimization-free: (1) Finetune-based: DreamBooth[3], CustomDiffusion[10], DreamMatcher[17]; (2) Encoder-based: ELITE[19], BLIP-Diffusion[20], IP-Adapter[43], MoMA[87], SSR-Encoder[88], Kosmos-G[86], Emu-2[89], $\lambda$-ECLIPSE[90], MS-Diffusion[91]. Among these methods, Kosmos-G[86] and Emu-2[89] are based on the multi-modal large language models.
Moreover, considering that there are also many methods focusing exclusively on human ID customized generation (a specific limited domain), we compare our approach with these methods in the human ID scenario to demonstrate our generalization, even though it is not the primary focus of this paper. We consider three recent state-of-the-art methods: PhotoMaker[23], Instant-ID[24], and PuLID[25].
@@ 4.1 Experimental Setups
TABLE II: Quantitative comparisons on multiple-subject customization. RealCustom++ achieves state-of-the-art performance: for text controllability, we achieve a 4.6% improvement on CLIP-B-T; for subject similarity, we achieve 6.34% and 3.9% improvements on CLIP-B-I and DINO-I, respectively.
@@ 4.2 Comparison with Stat-of-the-Arts
Quantitative main results. As shown in Table I, RealCustom++ outperforms existing methods across all metrics on both SD-v1.5 and SDXL: (1) For text controllability, on the more powerful base model SDXL, we achieve relative improvements of 7.01%, 7.48%, and 76.43% on CLIP-B-T, CLIP-L-T, and ImageReward, respectively. The significant improvement on ImageReward validates RealCustom++’s core motivation: non-conflict real-word subject representation and disentangling the influence scope between given texts and subjects. This approach generates customized images with much higher quality, as other subject-irrelevant regions are completely controlled by the given text without contamination from the subjects. (2) For subject similarity, we achieve state-of-the-art performance on CLIP-B-I, CLIP-L-I, and DINO-I, validating that RealCustom++ can generate more similar subjects in the subject-relevant regions. The figure of ”text-controllability (CLIP-B-T) versus subject-similarity (CLIP-B-I)” on SD-v1.5 in Figure 1(c) shows that the existing paradigm is trapped in the dual-optimum paradox, while RealCustom++ effectively eradicates it.
@@
Qualitative main results. As shown in Figure 9, RealCustom++ demonstrates superior zero-shot open-domain customization capability for a wide range of subjects, including humans, characters, buildings, animals, and uniquely shaped toys. It generates higher-quality custom images with better subject similarity and text controllability compared to existing methods. Furthermore, the disentangled influence scope between subjects and texts allows RealCustom++ to produce cleaner and more diverse backgrounds. For instance, in the second row, existing methods like Custom-Diffusion and MS-Diffusion inherit the pink background from the given subject image, while RealCustom++ generates a background that aligns more closely with the prompt ”on the beach”.
@@
TABLE III: Ablation study of using different Top-K $\gamma_{\text{scope}}$ ratios.
Figure 12: Illustration of gradually customizing the target real words into the given subjects for single-subject customization. The customized words are highlighted in red, with their attention maps gradually forming into the given subjects and details being added step by step. This process provides a more accurate image guidance mask for open-domain customization, while the remaining subject-irrelevant parts are completely controlled by the given text.
Comparison on multiple subjects customization. To further evaluate the multiple-subject customization ability, following the common practice of previous works[12, 91], we collect diverse subjects from previous works spanning various categories, including animals, characters, buildings, objects, etc., for a total of 30 different multiple subjects customized cases. The fully customized subjects and their prompts are provided in the Appendix. The subject similarity of multiple subjects is defined as the mean of each subject’s similarity. As illustrated in Table II, we outperform existing state-of-the-art works in multiple-subject customization: for text controllability, we achieve a 4.6% improvement on CLIP-B-T; for subject similarity, we achieve 6.34% and 3.9% improvements on CLIP-B-I and DINO-I, respectively. The qualitative comparison shown in Figure 10 demonstrates that RealCustom++ generates multiple-subject customization results with better text controllability and subject similarity than existing methods, aligning with the quantitative results in Table II. For example, in the second row, RealCustom++ successfully generates the scene of the given ”girl” playing chess with the given ”Stitch,” while existing methods struggle to keep both subjects consistent. Moreover, RealCustom++ handles diverse and complex interactions between different subjects, such as ”wearing” in the first row and ”playing guitar” in the fourth row. Both the quantitative and qualitative results further validate the superiority of our proposed new paradigm, RealCustom++.
No. Attention Setting CLIP-B-T $\uparrow$ CLIP-B-I $\uparrow$
(1) all resolution cross-attn 32.06 83.97
(2) low resolution cross-attn 32.20 84.23
(3) (2) + all resolution self-attn 31.08 85.22
(4) (2) + high resolution self-attn (Full Model) 33.43 87.32
TABLE IV: Ablation study on using different attentions to construct image mask guidance. The combination of low-resolution cross-attention and high-resolution self-attention provides the most accurate image mask for customization.
Comparison on human ID customization. Considering many existing customization methods focus solely on human ID customization[26, 24, 42, 25] (a specific limited domain), we further provide a qualitative comparison with these methods in Figure 11, demonstrating that our open-domain paradigm, RealCustom++, can also effectively generalize to this specific domain, even though human ID customization is not the primary focus of this paper. Specifically, RealCustom++ generates consistent IDs (as shown in the first two rows) while preserving elements such as clothing, hairstyles, and accessories (illustrated in the last three rows), which is challenging for existing methods.
Early Stop Step CLIP-B-T $\uparrow$ CLIP-B-I $\uparrow$
5 32.06 85.13
10 33.20 87.21
12 33.43 87.32
15 33.40 87.07
20 33.41 86.84
None (w/o early stop regularization) 33.41 86.67
TABLE V: Ablation study of mask calculation early stop regularization for more temporally stable image mask guidance.
@@ 4.3 Ablations
We ablate each component of our proposed RealCustom++ on SDXL, which mainly includes: (1) the adaptive mask guidance for disentangling the influence scope between the text and the subjects; (2) the cross-layer cross-scale projector for effectively encoding the subject image features; and (3) the curriculum training recipe for effective training.
Figure 13: Illustration of gradually customizing the native real words into the given subjects for multiple-subject customization. The different customized words in each case are highlighted in blue and green, respectively. We show that RealCustom++ provides accurate and decoupled guidance masks for each given subject, thereby achieving high-quality similarity for each subject.
@@
We then ablate the choice of using different attention to construct a more spatially accurate image mask guidance. As shown in Table IV, the combination of low-resolution cross-attention and high-resolution self-attention provides the most accurate image mask for customization, resulting in the best text controllability and subject similarity. We also visualize their differences in Figure 14, showing that (1) solely relying on cross-attention maps results in scattered mask guidance (e.g., the image guidance mask for the backpack, where holes exist in the center of the mask), causing the degradation of subject details; (2) using all-resolution self-attention maps makes the guidance mask over-focused, resulting in the degradation of mask accuracy.
The ablation study of early stop regularization for more temporally stable image mask guidance is provided in Table V. This aligns with our motivation shown in Figure 8, where the image guidance mask tends to converge in the middle diffusion steps, and reusing the middle step’s mask increases both similarity and controllability.
Figure 15: Visualization of different influence scope.
The qualitative and quantitative results of using different Top-K ratios $\gamma_{\text{scope}}$ are shown in Table III and Figure 15, respectively. We conclude that: (1) Within a proper range (experimentally, $\gamma_{\text{scope}}\in[0.15,0.4]$), the results are quite robust. (2) The maximum normalization in Eq. 15 is important for achieving both high similarity and controllability, as different regions in the selected parts have varying subject relevance and should be assigned different weights. (3) Too small or too large an influence scope will degrade similarity or controllability, respectively.
Finally, we ablate the impact of using different real words for customization in Figure 16. The customized real word is highlighted in red. We show that: (1) In the first row, regardless of how coarse-grained the word is, the customization results of RealCustom++ are quite robust. (2) In the second row, using completely different words to represent the given subject opens a door for a new application: novel concept creation. RealCustom++ combines these two concepts to create a new one, e.g., generating a tiger with the appearance of the given white Bichon Frise, as shown in the second row of Figure 16.
@@
TABLE VI: Ablation study of the cross-layer cross-scale projector.
Feature Combination Setting CLIP-B-T $\uparrow$ CLIP-B-I $\uparrow$
cross-scale element add + cross-layer token concat 29.94 86.33
cross-scale token concat + cross-layer token concat 33.65 85.62
TABLE VII: Ablation study of different feature combination.
Effectiveness of cross-layer cross-scale projector. As shown in Table VI, we ablate the design of our proposed cross-layer cross-scale projector, showing that both the cross-layer enhancement and the cross-scale enhancement bring substantial improvement over subject-similarity while cost little text-controllability, which validates our key idea of encoding subject images, i.e., keeping deep features’ ”dominating” position to avoid compromising their effectiveness of initial good alignment with the text condition
We next ablate the choice of different feature combinations in Table VII. The results validate our hypothesis that same-level features should be added element-wise, while different-level features should be concatenated token-wise to avoid conflicts.
@@
TABLE VIII: Ablation study of the curriculum training recipe.
Effectiveness of curriculum training recipe. As shown in Table VIII, we conclude that: (1) Training the model solely on <text-image> or <multiview> data significantly decreases either text controllability or subject similarity. This is because <text-image> data lacks subject pose diversity, while <multiview> data lacks open-domain subject diversity. (2) The curriculum data strategy achieves better CLIP-B-T and CLIP-B-I scores compared to naively using a 50% probability of <text-image> or <multiview> data. (3) The proposed curriculum cropping strategy further improves both metrics, highlighting the importance of promoting more stable initial convergence and better subsequent size generalization.
5 Discussion

### v2

@@ 3.3.2 Generation Branch
In the Generation Branch, $\boldsymbol{\hat{M}}$ is multiplied with the visual cross-attention results to mitigate any negative impacts on the controllability of the given texts in the subject-irrelevant regions. Specifically, Eq. 4 is reformulated as:
@@
where the necessary resize operation is applied to match the size of $\boldsymbol{\hat{M}}$ with the resolution of each cross-attention block. The denoised output of Generation Branch is denoted as $\boldsymbol{z_{t-1}}$. Classifier-free guidance [81] is applied to produce the next step’s denoised latent feature $\boldsymbol{z_{t-1}}$ as:
$\boldsymbol{z_{t-1}}=\epsilon_{\theta}(\varnothing)+\omega(\boldsymbol{z_{t-1}}-\epsilon_{\theta}(\varnothing)),$ (17)
where $\epsilon_{\theta}(\varnothing)$ is the unconditional denoised output and $\omega$ is the classifier-free guidance strength.
@@ 3.4 Extension To Multi-subjects Customization
In contrast to the existing pseudo-word paradigm, which learns the correspondence between each pseudo-word and its respective subject, RealCustom++ establishes a general alignment between vision conditions and all real words during training. This allows users to freely select different real words to customize various subjects during inference, making RealCustom++ inherently flexible and extendable to multiple-subject customization.
Extension to decouple multiple subjects from a single reference image (One2Many). Unlike previous pseudo word paradigm methods that require region-wise attention regularization or pre-defined masks to separate multiple subjects within a single reference, RealCustom++ can accomplish this task without any additional design: simply utilize different native real words to represent the target subjects in the reference images. As illustrated in Figure 2(b), using ’boy’ to customize only the target boy with green hair or simultaneously using ’boy’ and ’horse’ to customize both subjects, the disentanglement is naturally achieved thanks to the well-learned alignment between vision conditions and original text conditions.
Extension to compose multiple subjects from multiple reference images (Many2Many). Given $N$ reference images $\{x^{1},x^{2},...,x^{N}\}$, each of them represents a different given subject. We first use the proposed cross-layer cross-scale projector to encode them into $N$ vision conditions $\{f_{ci}^{1},f_{ci}^{2},...,f_{ci}^{N}\}$. Then the visual cross attention process in Eq.16 is repeatedly applied $N$ times for each vision conditions respectively, wherein the Top-K section process for the guidance mask is extended to a novel Multi-Subject Top-K selection as illustrated in Alg.1. The key of the Multi-Subject Top-K selection algorithm is to iteratively select the maximum scores for each subject while ensuring that these guidance masks do not overlap. After acquiring the selected attention map for each subject $\boldsymbol{\bar{M}}_{j},j\in[0,N-1]$, maximum normalization is applied to each:
$\boldsymbol{\hat{M}}_{j}=\frac{\boldsymbol{\bar{M}}_{j}}{\text{max}(\boldsymbol{\bar{M}}_{j})},j\in[0,N-1].$ (18)
@@
TABLE I: Quantitative comparisons. RealCustom++ outperforms existing works across all metrics on both SD-v1.5 and SDXL: (1) For text controllability, on the more powerful base model SDXL, we achieve improvements of 7.01%, 7.48%, and 76.43% on CLIP-B-T, CLIP-L-T, and ImageReward, respectively. The significant improvement on ImageReward validates that RealCustom++’s non-conflict subject representation and disentangled influence scope between texts and subjects generate higher-quality customized images (higher aesthetic score), as subject-irrelevant regions are controlled by the given text without contamination. (2) For subject similarity, we achieve state-of-the-art performance on CLIP-B-I, CLIP-L-I, and DINO-I, validating that RealCustom++ generates more similar subjects in subject-relevant regions. The evaluation benchmark is shown in Section 1.1 of the Appendix.
@@
TABLE II: Quantitative comparisons on multiple-subject customization. RealCustom++ achieves state-of-the-art performance: for text controllability, we achieve a 4.6% improvement on CLIP-B-T; for subject similarity, we achieve 6.34% and 3.9% improvements on CLIP-B-I and DINO-I, respectively. The evaluation benchmark is shown in Section 1.2 of the Appendix.
4 Experiments
4.1 Experimental Setups
Implementation Details. RealCustom++ is implemented based on the pretrained Stable Diffusion 1.5[8] and Stable Diffusion XL[9] to further validate the generalization. Stable Diffusion 1.5 [8] and Stable Diffusion XL [9] are two of the most widely used and authoritative pre-trained text-to-image base models for evaluating the performance of image generation algorithms. We use a mix of the generic <text-image> dataset (i.e., the filtered subset of Laion-5B [79] based on aesthetic score) and the <multi-view> dataset (i.e., MVImageNet[78]) as detailed in Sec.3.2.2. RealCustom++ is trained on 8 A100 GPUs for 160,000 iterations with a learning rate of 1e-4. For Stable Diffusion XL, we first train the model at a 512 resolution for 60,000 iterations and then further fine-tune it at a 1024 resolution for another 100,000 iterations. We use SigLip[77] and DINO[76] with a $384^{2}$ input image size as our image encoders, concatenating their last hidden states in the channel dimension to formulate the deep image features $\in\mathbb{R}^{729\times 2176}$. Therefore, the notation $n_{\text{image}}=729$ and $c_{\text{image}}=2176$. The shallow layer number $L=3$, and the selected shallow layers for SigLIP and DINO are {7, 13, 19} and {4, 10, 16}, respectively. These layers are also concatenated in the channel dimension correspondingly, resulting in triple-level shallow image features. Unless otherwise specified, for Stable Diffusion 1.5, the DDIM sampler [57] with 50 sample steps is used for sampling, and the classifier-free guidance $\omega$ is set to 12.5. For Stable Diffusion XL, the DDIM sampler [57] with 25 sample steps is used for sampling, and the classifier-free guidance $\omega$ is set to 7.5. The Top-K ratio $\gamma_{\text{scope}}$ is set to 0.2 by default. The timestep threshold $T_{\text{stop}}$ for mask calculation early stop regularization is set to 25 for SD-v1.5 and 12 for SDXL.
(1) Subject-Similarity: We use the state-of-the-art segmentation model, SAM [88], to segment the subject, and then evaluate with both CLIP-I and DINO [76] scores. These scores represent the average pairwise cosine similarity of CLIP or DINO embeddings of the segmented subjects in generated and given images. For a more robust evaluation, we adopt two variants of CLIP: CLIP ViT-B/32 (denoted as CLIP-B-I) and CLIP ViT-L/14 (denoted as CLIP-L-I). Since the larger CLIP ViT-L/14 variant represents images with more fine-grained features, it can therefore reflect more fine-grained subject similarity.
Figure 9: Qualitative comparison with existing state-of-the-art customization methods. RealCustom++ produces higher quality results with better subject similarity and text controllability compared to existing methods. Additionally, RealCustom++ demonstrates superior subject diversity in poses and expressions (e.g., ”smile on her face” in the first row). Furthermore, RealCustom++ generates customization results with cleaner and more diverse backgrounds. For instance, in the second row, existing methods like Custom-Diffusion and MS-Diffusion inherit the pink background from the given subject image, while RealCustom++ generates a background that aligns more closely with the prompt ”on the beach”.
Evaluation Benchmarks. Following previous works, we use the prompt ”a photo of [category]” to generate images for calculating similarity. The complete set of subject images and editing prompts used for evaluation are detailed in Section 1.1 of the Appendix. This section includes 50 different subjects, such as daily necessities, buildings, animals, cartoon characters, and humans, as well as over 50 different types of editing text prompts, including re-contextualization, shape alteration, and outfit changes, resulting in a total of 740 cases for a thorough evaluation. These editing prompts and subject images are derived from the standard DreamBench[3] and are further supplemented with more challenging open domain cases provided by users in real-world applications, such as cartoon characters and buildings, for a more comprehensive evaluation than our conference version.
Prior SOTAs. We compare with existing state-of-the-art text-to-image customization methods includes both finetune-based and encoder-based: (1) Finetune-based: DreamBooth[3], CustomDiffusion[10], DreamMatcher[17]; (2) Encoder-based: ELITE[19], BLIP-Diffusion[20], IP-Adapter[43], MoMA[83], SSR-Encoder[84], Kosmos-G[82], Emu-2[85], $\lambda$-ECLIPSE[86], MS-Diffusion[87]. Among these methods, Kosmos-G[82] and Emu-2[85] are based on the multi-modal large language models.
@@ 4.2 Comparison with Stat-of-the-Arts
Quantitative main results. As shown in Table I, RealCustom++ outperforms existing methods across all metrics on both SD-v1.5 and SDXL: (1) For text controllability, on the more powerful base model SDXL, we achieve relative improvements of 7.01%, 7.48%, and 76.43% on CLIP-B-T, CLIP-L-T, and ImageReward, respectively. The significant improvement on ImageReward validates RealCustom++’s core motivation: non-conflict real-word subject representation and disentangling the influence scope between given texts and subjects. This approach generates customized images with much higher quality, as other subject-irrelevant regions are completely controlled by the given text without contamination from the subjects. (2) For subject similarity, we achieve state-of-the-art performance on CLIP-B-I, CLIP-L-I, and DINO-I, validating that RealCustom++ can generate more similar subjects in the subject-relevant regions. The figure of ”text-controllability (CLIP-B-T) versus subject-similarity (CLIP-B-I)” on SD-v1.5 in Figure 1(c) shows that the existing paradigm is trapped in the dual-optimum paradox, while RealCustom++ effectively eradicates it. (3) Compared with our conference version[31], we achieve 6.06% improvement on text controllability (i.e., CLIP-L-T) and 2.31% improvement on subject similarity (i.e., DINO-I), which validates the effectiveness of our key algorithm redesign. Moreover, we also achieve 56.6% improvement to our conference version on the image quality.
@@
Qualitative main results. As shown in Figure 9, RealCustom++ demonstrates superior zero-shot open-domain customization for various subjects. It generates higher-quality custom images with better subject similarity and text controllability compared to existing methods. Additionally, the disentangled influence scope between subjects and texts allows RealCustom++ to produce cleaner and more diverse backgrounds. For instance, in the second row, existing methods like Custom-Diffusion and MS-Diffusion inherit the pink background from the given subject image, while RealCustom++ generates a background that aligns more closely with the prompt ”on the beach”.
Comparison on multiple subjects customization. To further evaluate the multi-subject customization ability, we follow the common practice of previous works [12, 87] by collecting diverse subjects from prior studies. These subjects span various categories, including animals, characters, buildings, and objects, resulting in a total of 30 customization cases. The fully customized subjects and their prompts are provided in Section 1.2 of the Appendix. The subject similarity of multiple subjects is defined as the mean of each subject’s similarity. As illustrated in Table II, we outperform existing state-of-the-art works in multiple-subject customization: for text controllability, we achieve a 4.6% improvement on CLIP-B-T; for subject similarity, we achieve 6.34% and 3.9% improvements on CLIP-B-I and DINO-I, respectively. The qualitative comparison shown in Figure 10 demonstrates that RealCustom++ generates multiple-subject customization results with better text controllability and subject similarity than existing methods, aligning with the quantitative results in Table II. For example, in the second row, RealCustom++ successfully generates the scene of the given ”girl” playing chess with the given ”Stitch,” while existing methods struggle to keep both subjects consistent. Moreover, RealCustom++ handles diverse and complex interactions between different subjects, such as ”wearing” in the first row and ”playing guitar” in the fourth row. Both the quantitative and qualitative results further validate the superiority of our proposed new paradigm, RealCustom++.
Comparison on human ID customization. The aforementioned results on open-domain subjects demonstrate our superior customization performance. Considering that many concurrent works focus exclusively on human ID customization (a specific domain), we also provide a comparison in Section 2.1 of the Appendix to further demonstrate the strong generalization of RealCustom++ in this domain. The results show that RealCustom++ generates more consistent human IDs with natural poses compared to existing methods. Additionally, RealCustom++ effectively preserves clothing, hairstyles, and accessories, which is crucial for real-world applications like film production and remains challenging for current methods.
No. Attention Setting CLIP-B-T $\uparrow$ CLIP-B-I $\uparrow$
(1) all resolution cross-attn 32.06 83.97
(2) low resolution cross-attn 32.20 84.23
(3) (2) + all resolution self-attn 31.08 85.22
(4) (2) + high resolution self-attn (Full Model) 33.43 87.32
TABLE III: Ablation study on using different attentions to construct image mask guidance. The combination of low-resolution cross-attention and high-resolution self-attention provides the most accurate image mask for customization.
@@
TABLE IV: Ablation study of using different Top-K $\gamma_{\text{scope}}$ ratios.
@@ 4.3 Ablations
We ablate each key component of RealCustom++ on SDXL.
@@
We then ablate the choice of using different attention to construct a more spatially accurate image mask guidance. As shown in Table III, the combination of low-resolution cross-attention and high-resolution self-attention provides the most accurate image mask for customization, resulting in the best text controllability and subject similarity. The visualization could be found be Section 2.2 of the Appendix.
Figure 11: Illustration of customizing real words into given subjects for single-subject customization. The customized words are highlighted in red, with their attention maps gradually forming the subjects’ shape and details step by step. This process provides accurate image guidance masks for open-domain subjects, while subject-irrelevant parts are controlled by the given text.
Figure 12: Illustration of customizing real words into given subjects for multiple-subject customization. The customized words are highlighted in blue and green, respectively. RealCustom++ provides accurate and decoupled guidance masks for each subject.
The qualitative results of using different Top-K ratios $\gamma_{\text{scope}}$ are shown in Table IV. We conclude that: (1) Within a proper range (experimentally, $\gamma_{\text{scope}}\in[0.15,0.4]$), the results are quite robust. (2) The maximum normalization in Eq. 15 is important for achieving both high similarity and controllability, as different regions in the selected parts have varying subject relevance and should be assigned different weights. (3) Too small or too large an influence scope will degrade similarity or controllability, respectively.
@@
TABLE V: Ablation study of the cross-layer cross-scale projector.
Effectiveness of cross-layer cross-scale projector. As shown in Table V, we ablate the design of our proposed cross-layer cross-scale projector, showing that both the cross-layer enhancement and the cross-scale enhancement bring substantial improvement over subject-similarity while cost little text-controllability, which validates our key idea of encoding subject images, i.e., keeping deep features’ ”dominating” position to avoid compromising their effectiveness of initial good alignment with the text condition
@@
TABLE VI: Ablation study of the curriculum training recipe.
Effectiveness of curriculum training recipe. As shown in Table VI, we conclude: (1) Training solely on <text-image> or <multiview> data significantly decreases either text controllability or subject similarity. <text-image> data lacks subject pose diversity, while <multiview> data lacks open-domain subject diversity. (2) The curriculum data strategy achieves better CLIP-B-T and CLIP-B-I scores compared to using a 50% probability of <text-image> or <multiview> data. (3) The proposed curriculum cropping strategy further improves both metrics, highlighting the importance of stable initial convergence and better subsequent size generalization.
Due to limited space, more ablation results can be found in the appendix: (1) The ablation study of early stop regularization in Section 2.3 of the Appendix, validating its effectiveness for temporally stable image mask guidance. (2) The visualization of qualitative results using different Top-K ratios $\gamma_{\text{scope}}$ in Section 2.4 of the Appendix, validating our quantitative results in Table IV. (3) The ablation of the impact of using different real words for customization in Section 2.5 of the Appendix, validating the robustness of RealCustom++. (4) The ablation of using different feature combinations in Section 2.6 of the Appendix, validating our hypothesis that same-level features should be added element-wise, while different-level features should be concatenated token-wise to avoid conflicts. (5) The limitations of RealCustom++ are discussed in detail in Section 3 of the Appendix, where the main limitation lies in that if the pre-trained text-to-image models cannot generate the concept of the target real word, RealCustom++ cannot customize the specific subject generation of this target word.